OpenStudy (anonymous):

A parabola has a vertex at (-1, 0) and opens down. What is the equation of the parabola? y = -x^2 - 1 y = -(x - 1)^2 y = -(x + 1)^2

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

last one?

Parth (parthkohli):

There's not enough information in the question, because the general vertex-form a parabola is\[f(x) = a(x-h)^2 + k\]where \((h,k)\) is the vertex. All of your choices have \(a=-1\), so might as well assume that.\[f(x) = -1\cdot(x-h)^2 + k\]It is given that \(h = -1\) and \(k =0\).

OpenStudy (perl):

if you use f(x) = a(x-h)^2 + k , you are given h = -1, and k = 0. plug that in. also you know that a<0 since the parabola points downwards

OpenStudy (anonymous):

thanks